OpenStudy (anonymous):

A coffee pot makes 6 cups of coffee in 1 1/2min. How many cups of coffee can the coffee pot make in 1 min? cups

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[1\frac{ 1 }{ 2 }=1+\frac{ 1 }{ 2 }=\frac{ 2*1+1 }{ 2 }=\frac{ 3 }{ 2 }\] \[no.~of~ cups~ per~ minute=6\div \frac{ 3 }{ 2 }=6\times \frac{ 2 }{ 3 }=?\]

OpenStudy (anonymous):

4

OpenStudy (anonymous):

correct